OpenStudy (anonymous):

Find all complex solutions of 4x^(2)-5x+2=0

jhonyy9 (jhonyy9):

discriminant = ?

jhonyy9 (jhonyy9):

D=25-32=-7

jhonyy9 (jhonyy9):

x_1,2=(5+/- sqrt(-7))/8 =(5+/-isqrt7)/8

OpenStudy (anonymous):

i have no idea if you tryung to find the discrimant

jhonyy9 (jhonyy9):

formula of discriminant D=b^2 -4ac

OpenStudy (anonymous):

probably not since its only asking to find al complex solutions

OpenStudy (anonymous):

what is your final answer?

jhonyy9 (jhonyy9):

there are the complex solutions - so because the discriminant is less than zero so than the roots are complex numbers

jhonyy9 (jhonyy9):

x_1,2=(5+/-isqrt7)/8

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

so x=(5+/-isqrt7)/8???

jhonyy9 (jhonyy9):

yes

OpenStudy (anonymous):

or are 1 and 2 part of your solution

jhonyy9 (jhonyy9):

not is part just hence in this way note those roots x_1 and x_2

OpenStudy (anonymous):

thanks are those slashes divide simbols? or is that how you seperate yo numbers?

jhonyy9 (jhonyy9):

this is just to separate these two roots like x_1 and x_2 ok ?

OpenStudy (anonymous):

well can you write the answer without the slashes(if you are seperating them) so im not confused

jhonyy9 (jhonyy9):

yes just write it without ,,_" ok ? like x1,2 =(5+/- isqrt7)/8 or x_1=(5-isqrt7)/8 x_2=(5+isqrt7)/8

OpenStudy (anonymous):

thank you and God bless america.

jhonyy9 (jhonyy9):

good luck yw bye